OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = –2 and a5 = –512? –8,192 2,048 –2,048 8,192

OpenStudy (lalaly):

\[a_n = a_1 \times r^{(n - 1)}\]we have a1 and a5 so solve for r\[- 512 = (- 2) \times r^{(5 - 1)}\]\[r^4=256\]\[r=4\] now \[a_7=a_1 \times r^{7-1}\]\[a_7= -2 \times 4^6\]

OpenStudy (lalaly):

the answer is the first one

OpenStudy (anonymous):

THank you soo much :)

OpenStudy (lalaly):

your welcome :D